Analyze the annual reports of Flour Mills of Fiji (2020), RB Patel Ltd (2020) and Paradise Beverages Ltd (2019) from South Pacific Stock Exchange (SPSE) and answer the following questions. 1. Calculate the share value using Dividend Growth Model, assuming dividend will grow 2% for Flour Mills of Fiji, 5% for RB Patel Ltd and 8% for Paradise Beverages constantly for the companies on their respective base year. 2. Calculate the Earnings Per Share (EPS) and the Market Capitalization for each of the companies. 3. Comment on the share value of the companies given the Dividend Growth Model (Question 1) and EPS (Question 2)


1
Expert's answer
2021-05-24T08:59:30-0400

1.solving by the Gordon model:

"P=\frac{D(1+q)}{i-q}"

let the dividends be 5, 6,7

discount rate 10%

"P1=\frac{D(1+q)}{i-q}=\frac{5(1+0,02)}{0,1-0,02}=63,75"

"P2=\frac{D(1+q)}{i-q}=\frac{6(1+0,05)}{0,1-0,05}=105"

"P3=\frac{D(1+q)}{i-q}=\frac{7(1+0,08)}{0,1-0,08}=270"

2.Let the profit be 2000,5000,6000, the number of shares is 2000, 5000, 6000, then

"EPS=\frac{P}{number of shares}"

"EPS1=\frac{2000}{2000}=1"


"EPS2=\frac{5000}{5000}=1"


"EPS3=\frac{6000}{6000}=1"

the Market Capitalization

"МС=P\times number of shares"

"МС1=63.75\times 2000=127 500"

"МС2=105\times 5000=525 000"

"МС3=270\times 6000=1 620 000"

3.the bigger the dividend, the bigger the price, the bigger the capitalization